Three students, A, B and C, received a Stadio Mathematics award for their top performance. They have decided to split the award as follows: i. Student A receives R10 000 plus 1 5 of what then remains; ii. Student B then receives R16 000 plus 1 4 of what then remains; and iii. Student C then receives the rest, which is R18 000. How much is the original monetary award? Which student receives the most money?


1
Expert's answer
2021-05-07T09:54:00-0400

Let original monetary award be x

Student A receives 10,000 and "\frac{1}{5}" of remaining

"A=10,000+\frac{1}{5}(X-10,000)........(i)"

Student B receives 16,000 and "\frac{1}{4}" of remaining

"B=16,000+\frac{1}{4}(X-16,000-A).....(ii)"

Student C receives the rest which is 18000

"X-A-B=18,000.......(iii)"

From (i)(ii)(iii)

"=X-[10,000+\frac{1}{5}(X-10,000)]-[16,000+\frac{1}{4}(X-16,000-A)]=18000"

"=X-8000-\frac{1}{5}X-12000-\frac{1}{4}X+\frac{A}{4}=18000"

"=X-8000-\frac{1}{5}X-12000-\frac{1}{4}X+\frac{1}{4}(8000+\frac{1}{5}X)=18000"


"=\frac{12}{20}X=18000+18000"

"X=60,000"

"R 60,000" is the original monetary award.


Student A "=10,000+\frac{1}{5}×(60000-10000)"

"=R 20000"

Student B

"=16000+\frac{1}{4}×(60000-16000-20000)"

"=R 22000"

Student C

"=R 18000"

Student B receives the most money.